# UniCUE: Unified Recognition and Generation Framework for Chinese Cued Speech Video-to-Speech Generation

### v1

@@ 3.2. Pose-aware Visual Processor
Considering the spatiotemporal correlation between hand coding, lip reading, and their semantic content, both CSV2S and CSR require precise localization of lip and hand keypoints. This necessity motivates the visual encoder to capture fine-grained pose features. To address this, we develop a pose-aware visual processor to construct mixed visual representations by synergistically integrating pose and video embeddings, as shown in Figure 3.
Specifically, given the video input and pose input , a tokenizer first splits them into patch tokens. Let the ViT-based image encoder be denoted as . We extract the video features and pose features from the given video tokens and pose tokens , respectively:
where and . Here, represents the number of visual patch tokens and denotes the dimension of visual embedding. The video features and pose features are then concatenated along channel dimension, apply a trainable projector layer to form the mixed visual embedding :
where MLP refers to the projection layer. The CSR path supervises the training of the pose-aware visual processor, guiding it to adapt to cuer-specific CS expressions and learn fine-grained semantic representations.
@@ 3.3. Semantic Alignment Pool
To further enhance fine-grained semantic consistency between visual representation and linguistic content, we propose a semantic alignment mechanism that synchronizes visual, pose, and textual modalities through contrastive learning. Specifically, a ViT-based text encoder first projects tokens of ground-truth transcripts into embeddings , sharing the same latent space with visual and pose features extracted by the pose-aware processor. We formulate the alignment objective as a multi-modal embedding alignment Loss:
where Cos(,) denotes the cosine similarity. The total semantic alignment loss is calculated via:
By individually aligning video and pose features with text embeddings, the model can effectively integrate complementary cues, enhancing its ability to understand visual cues for speech generation.
@@ 3.4. VisioPhonetic Adapter
The VisioPhonetic adapter (VPA) acts as a critical translator between the learned pose-aware visual processor in CSR and the diffusion-based speech generator in CSV2S, addressing the representation gap caused by their heterogeneous architectures. As illustrated in Figure 4, this lightweight module employs a sequential architecture to progressively refine visual-semantic representations into diffusion-compatible conditioning signal:
which includes two MLPs and a Q-Former-style (Li et al., 2023) cross-attention layer. The cross-attention mechanism is specifically designed to bridge the representation gap through learnable semantic queries. Formally, given the mixed visual embedding , the process is formulated as: where denotes the learnable query vector initialized from the diffusion model’s latent space, and , , are projection matrices.
The VPA effectively aligns visual information with the semantic requirements of speech generation, further mapping the semantic connections between hand-lip dynamics, phonetic constraints, and speech. temporally related information from the visual features, thereby improving the accuracy, synchronization, and cuer-specific expression in speech.
3.5. Training and Inference
The training process is divided into two sequential phases to ensure effective knowledge transfer from visual understanding to speech synthesis. Stage 1: CSR focuses on learning fine-grained spatiotemporal representations by jointly optimizing the pose-aware visual processor, the semantic alignment pool, text encoder and decoder. This stage trains the model using the cross-entropy loss (as defined in Eq. 1) and the contrastive alignment loss (as defined in Eq. 7), formulated as: Stage 2: CSV2S leverages the learned visual representations from Stage 1 to train the VPA network and diffusion model with the objective (denoted in Eq. 3).
During inference, the framework operates in a streamlined single-path mode to synthesize speech directly from cued CS videos. As illustrated in Figure 5 (b), given an CS video and its corresponding pose, the trained pose-aware visual processor extracts the visual mixture representation , which encodes both semantic content and cuer characteristics. A noise sample , sampled from a standard Gaussian distribution , is fed into the trained LDM to predict the speech latent code conditioned on the visual mixture representation over timesteps:
Subsequently, the predicted speech latent code is decoded into a speech mel-spectrogram via VAE decoder, and a pre-trained Vocoder (Liu et al., 2023) is used to convert the mel-spectrogram into a speech waveform. Through this series of steps, the CSV2S system is able to generate natural speech that is consistent with the cuer’s characteristics based on the input CS video and its corresponding pose information.
| Dataset | Cuers | Sentences | Resolution | FPS |
| French CS (Liu et al., 2018) | 1 normal-hearing cuer | 238 | 50 | |
| British CS (Liu et al., 2019) | 1 normal-hearing cuer | 97 | 25 | |
| Chinese CS (Lei et al., 2024b) | 4 normal-hearing cuers | 4000 | 30 | |
| Ours (Chinese CS) | 6 normal-hearing and 8 hearing-impaired cuers | 11282 | 30 | |
| Method | Normal-hearing cuers | Hearing-impaired cuers | |||||
| WER | LSE-D | LSE-C | FAD | WER | LSE-D | LSE-C | |
| GT | - | 7.250 | 0.200 | - | - | - | - |
| CMML (Liu and Liu, 2023)+TTS | 0.614 | 9.327 | 0.124 | 4.908 | 1.102 | 9.825 | 0.090 |
| CNN+FLASH (Hua et al., 2022)+TTS | 0.956 | 9.773 | 0.090 | 5.747 | 0.998 | 9.943 | 0.070 |
| CNN+Linformer (Wang et al., 2020)+TTS | 0.991 | 9.874 | 0.082 | 5.944 | 1.092 | 10.014 | 0.067 |
| CNN+Performer (Choromanski et al., 2020)+TTS | 0.945 | 9.873 | 0.105 | 5.429 | 0.988 | 9.876 | 0.079 |
| CNN+Cosformer (Zhen et al., 2022)+TTS | 1.069 | 10.063 | 0.076 | 6.355 | 1.229 | 10.115 | 0.064 |
| EcoCued (Liu et al., 2024)+TTS | 0.672 | 9.495 | 0.114 | 5.326 | 1.621 | 9.846 | 0.087 |
| Ours (CSR)+TTS | 0.030 | 8.461 | 0.158 | 3.653 | 0.254 | 9.045 | 0.127 |
| Ours (UniCUE) | 0.056 | 7.432 | 0.194 | 1.749 | 0.293 | 8.716 | 0.153 |
@@ 4. Experiment
4.1. Dataset
To further enhance the model’s generalization ability and validate its practical effectiveness in facilitating communication for hearing-impaired individuals, we establish a new Chinese CS dataset by incorporating CS data from 2 normal-hearing and 8 hearing-impaired cuers with 4 normal-hearing cuers data in (Lei et al., 2024b), increasing the total number of sentences in the dataset to 11,282. A comparison between our dataset and existing CS datasets is presented in Table 1. Specifically, we invited 2 normal-hearing cuers and 8 hearing-impaired cuers to simultaneously utter and encode a set of text sentences provided by (Lei et al., 2024b). The 10 cuers underwent systematic training to ensure they could perform Mandarin Chinese CS fluently and accurately. It is important to note that our dataset was collected with the explicit consent of all cuers and is available for open-source use. We use CS data from 6 normal-hearing cuers to train the model, splitting this data into training and test sets in a ratio to ensure effective model training and preliminary validation. To more comprehensively evaluate the practical effectiveness in facilitating communication for hearing-impaired individuals, we included all the CS data from 8 hearing-impaired cuers in the test set.
4.2. Implementation Details
Architecture Details. To construct the CSV2S path, we employ a speech generation pipeline consisting of a VAE encoder and decoder, an LDM, and a Vocoder from AudioLDM (Liu et al., 2023). We initialize the pipeline by loading the pre-trained weights of the VAE and LDM. For the CSR path, we utilize a conditional translation LLM, MBart (Liu, 2020), which comprises a tokenizer, an image encoder, a text encoder, and a text decoder. The pipeline is initialized by loading the pre-trained weights with multi-language corpus222https://huggingface.co/facebook/mbart-large-50.
Training Details. We extract CS pose information—including hands, face, and body—from videos using OpenPose333https://github.com/CMU-Perceptual-Computing-Lab/openpose. Raw video frames and pose images are resized to , while the audio is resampled to 16 kHz and converted into mel-spectrograms. The entire learning framework is implemented in PyTorch, and all experiments are conducted on an NVIDIA RTX A6000. We use the Adam optimizer with a learning rate of for end-to-end training, setting the mini-batch size to 2.
Comparison Methods. To validate the effectiveness of our framework, we conduct extensive comparisons between our method UniCUE (CSV2S with CSR-transferred information) and the following methods: 1) our single CSV2S: an end-to-end trained CSV2S without the shared CSR information; 2) our CSR: comprising of the pose-aware visual processor and text decoder to output CS linguistic text (shown in Figure 3); 3) SOTA CSR methods: including CMML (Liu and Liu, 2023), EcoCued (Liu et al., 2024), and various combinations of CNNs and transformers, such as FLASH (Hua et al., 2022), Linformer (Wang et al., 2020), Performer (Choromanski et al., 2020), and Cosformer (Zhen et al., 2022). For fairness, we reproduce these CSR methods on our dataset using the same experimental setup described in their respective papers. Since CSR methods only output phoneme-level transcriptions, we propose a two-stage conversion pipeline to synthesis speech, as illustrated in Figure 5 (a). The process consists of: 1) converting CSR outputs into Chinese pinyin sequences using rule-based CS patterns, followed by translating the pinyin into Chinese characters with QWen toolkit444https://github.com/QwenLM/Qwen;2) generating the final waveforms using SOTA TTS model F5-TTS (Chen et al., 2024) with reference speech.
| Method | Normal-hearing cuers | Hearing-impaired cuers | |||||
| WER | LSE-D | LSE-C | FAD | WER | LSE-D | LSE-C | |
| GT | - | 7.250 | 0.200 | - | - | - | - |
| CSR††+TTS | 0.125 | 8.847 | 0.129 | 4.231 | 0.307 | 9.672 | 0.098 |
| CSR‡+TTS | 0.097 | 8.716 | 0.135 | 4.063 | 0.291 | 9.472 | 0.106 |
| CSR+TTS | 0.030 | 8.461 | 0.158 | 3.653 | 0.254 | 9.045 | 0.127 |
| CSV2S† | 0.288 | 8.379 | 0.141 | 3.205 | 0.403 | 8.984 | 9.128 |
| CSV2S | 0.258 | 8.106 | 0.147 | 2.864 | 0.379 | 8.912 | 0.130 |
| UniCUE* | 0.104 | 7.741 | 9.172 | 2.219 | 0.332 | 8.884 | 0.139 |
| UniCUE | 0.056 | 7.432 | 0.194 | 1.749 | 0.293 | 8.716 | 0.153 |
4.3. Evaluation Metrics
To evaluate the performance of the synthesized speech, we employ three categories of metrics focusing on linguistic accuracy, audio-visual synchronization, and speech quality. For linguistic accuracy, we measure the Word Error Rate (WER) between the recognized text of generated speech and ground truth transcripts. The transcripts are normalized by removing punctuation and converting numbers to words prior to alignment. To measure the audio-visual synchronization between the generated speech and the corresponding video, we utilize SyncNet (Chung and Zisserman, 2016) and report two key metrics: LSE-D (the temporal distance between audio and video) and LSE-C (the confidence score). These values are computed using the pre-trained SyncNet model555https://github.com/joonson/syncnet-python. For speech quality, Fréchet Audio Distance (FAD) (Kilgour et al., 2018) is employed to assess the naturalness of generated speech by measuring the distribution divergence between real and synthesized audio in a pretrained VGGish embedding space (Hershey et al., 2017).
@@ 4.4. Comparison with SOTA Methods
We conduct comparison with SOTA methods. As shown in Table 2, our framework demonstrates superior performance: 1) Thanks to the pose-aware visual processor and semantic alignment pool, our CSR model achieves WER scores of 0.030 for normal-hearing cuers and 0.254 for hearing-impaired cuers, significantly outperforming previous SOTA CSR methods; 2) Leveraging the strong understanding capabilities of CSR, our UniCUE model also achieves lower WER scores. Additionally, our UniCUE surpasses others in FAD, LSE-D, and LSE-C metrics, showcasing high-quality, temporally aligned, and expressive speech generation.
@@ 4.5. Ablation Studies
To systematically validate the contribution of each component, we conduct controlled ablation studies on both normal-hearing and hearing-impaired test data. The results are listed in Table 3.
The key findings are analyzed as follows: 1) Unified Training Paradigm. Compared with the single CSV2S baseline, our UniCUE reduces WER by 62% (0.056 vs. 0.258) for normal-hearing cuers and 22.7% (0.293 vs. 0.379) for hearing-impaired groups. This verifies that transferring fine-grained semantic knowledge from CSR to CSV2S through shared visual representation learning effectively alleviates the challenge of complex semantic correlation.
2) Visual Processor Design. We analyze the impact of pose-aware visual processor by comparing CSR‡+TTS vs. CSR+TTS and CSV2S† vs. CSV2S. Models relying solely on raw video features exhibit suboptimal performance due to their limited capacity to capture fine-grained motion patterns. This suggests that unprocessed video frames contain substantial interference from redundant visual information. By integrating pose estimation features as supplementary inputs, the model gains enhanced fine-grained awareness of motions, leading to significant improvements in both CSR and CSV2S tasks. Furthermore, the pose features enhances model robustness against inter-cuer variability. As illustrated in Figure 6, our model maintains consistent performance across six normal-hearing cuers when speaking same linguistic content, demonstrating the importance of introducing pose-aware information. 3) Semantic Alignment Mechanism. Disabling the semantic alignment pool (CSR††+TTS) weakens the pose-text correlation, causing 17.9% relative WER increase (0.030 vs. 0.125). The contrastive alignment loss enforces spatiotemporal consistency between visual cues and CS phonemes, which is particularly beneficial for accurate semantic content. 4) VPA Mechanism. Removing the VPA (UniCUE*) results in an 85.7% degradation in FAD (1.749 vs. 2.219), indicating that the cross-attention-based VPA effectively bridges the representation gap by adaptively selecting critical visual-semantic cues relevant to speech characteristics through learnable query interaction. 5) Qualitative Analysis. To more comprehensively demonstrate the advantages of our unified framework UniCUE, we also provide qualitative comparison of CSR+TTS, CSV2S, and UniCUE. As shown in Figure 7, UniCUE achieves more accurate temporal alignment with the ground truth, indicating its superior ability to capture temporal dynamics within the visual modality. In contrast, CSR and CSV2S exhibit noticeable temporal drift.

### v3

@@ 3.2. Pose-aware Visual Processor
Considering the strong spatiotemporal correlation between hand coding, lip movement, and their underlying semantic content, both CSV2S and CSR require accurate modeling of lip and hand motion patterns. This necessitates a visual encoder capable of capturing fine-grained and temporally coherent features. While video frames offer rich appearance information, they often suffer from redundancy and visual ambiguity. In contrast, pose maps provide a compact, structured, and noise-resilient representation of motion dynamics. To leverage the complementary strengths of both modalities, we design a pose-aware visual processor that constructs fused visual representations, as shown in Figure 3.
Specifically, the input to the processor consists of video frames and pose maps , both formatted as tensors of shape , where indicates the frame lengths, and denotes the spatial resolution. The processor comprises two main components. First, a shared visual encoder extracts spatiotemporal features from both modalities via a sequential architecture: a 2D ResNet backbone extracts frame-wise spatial features, which are stacked along the temporal axis and passed through a 1D temporal convolution to model short-term motion patterns. The resulting sequence is then fed into a Transformer encoder to capture long-range temporal dependencies across frames. This process yields the video features and pose features , where , with denoting the embedding dimension and being the total number of tokens, where is the number of spatial patches per frame. Second, the projection layer integrates the two feature streams. The video and pose features are concatenated along the channel dimension and passed through a multi-layer perceptron (MLP), consisting of two linear layers with ReLU activation and LayerNorm, to produce the final mixed visual representation:
The fused representation serves as a unified visual embedding that drives both recognition and generation pathways. In the subsequent modules, this representation is semantically aligned with linguistic content and refined for diffusion-based speech synthesis.
@@ 3.3. Semantic Alignment Pool
To further enhance semantic consistency between visual representation and linguistic content, we introduce a semantic alignment mechanism that aligns video, pose, and textual modalities through contrastive learning. Specifically, a ViT-based text encoder encodes the ground-truth transcript tokens into text embeddings . The visual features and pose features , extracted by the pose-aware visual processor, are projected into a shared latent space via learnable linear layers. The text embedding is similarly projected. We adopt a contrastive loss across the batch, treating each video-text and pose-text pair from the same sample as a positive pair, and all others as negatives. The loss is denoted as:
where denotes the cosine similarity between normalized embeddings. The total semantic alignment loss is calculated as:
By enforcing this high-level alignment, the model is encouraged to extract complementary and discriminative semantics from visual modalities.These aligned features not only enhance linguistic recognition in CSR, but also offer semantically grounded condition for accurate speech synthesis.
@@ 3.4. VisioPhonetic Adapter
While the CSR-derived embeddings capture rich visual-linguistic semantics, they remain mismatched in format and granularity for direct use in diffusion-based speech generation. To bridge this modality gap, we propose the VisioPhonetic Adapter (VPA), which transforms semantically aligned visual features into a phonetic-aware conditioning signal suitable for the LDM. As illustrated in Figure 4, this lightweight module employs a sequential architecture to progressively refine visual-semantic representations into a diffusion-compatible conditioning signal:
which includes two MLPs and a Q-Former-style (Li et al., 2023) cross-attention layer. We use learnable semantic queries , which is initialized by computing the average latent representation from ground-truth mel-spectrograms encoded by the pretrained VAE. This provides a phonetic-aware initialization aligned with the diffusion model’s target space. These queries act as phonetic slots to extract and reorganize relevant patterns from . The cross-attention mechanism operates as: The adapted features serve as the final interface between visual understanding and speech synthesis, ensuring that the generated audio is not only temporally coherent but also linguistically faithful to the video input.
Dataset | Cuers | Sentences | Character | Word | Resolution | FPS | |
French CS (Liu et al., 2018) | 1-H | 238 | 12872 | - | 50 | ||
British CS (Liu et al., 2019) | 1-H | 97 | 2741 | - | 25 | ||
MCCS (Lei et al., 2024b) | 4-H | 4000 | 131608 | 42256 | 30 | ||
Unified-HI (Ours) |
11282 | 350333 | 112664 | 30 | ||
Method | Normal-hearing cuers | Hearing-impaired cuers | |||||||
WER | LSE-C | LSE-D | DNSMOS | STOI | WER | LSE-C | LSE-D | DNSMOS | |
GT | - | 7.274 | 7.314 | 2.79 | - | - | - | - | - |
CMML | 0.663 | 4.135 | 9.241 | 1.24 | 0.11 | 0.924 | 2.141 | 10.132 | 1.03 |
EcoCued | 0.657 | 4.327 | 9.146 | 1.28 | 0.12 | 0.917 | 2.165 | 10.079 | 1.07 |
CSR (Ours) | 0.186 | 4.874 | 9.125 | 2.53 | 0.57 | 0.224 | 3.342 | 9.315 | 2.29 |
Lip2Speech | 0.803 | 4.215 | 9.367 | 1.03 | 0.05 | 0.989 | 2.424 | 10.816 | 0.02 |
LipVoicer | 0.754 | 4.361 | 9.226 | 1.12 | 0.08 | 0.971 | 2.623 | 10.517 | 0.04 |
CSV2S (Ours) | 0.374 | 6.245 | 7.962 | 2.27 | 0.42 | 0.422 | 5.938 | 8.347 | 2.04 |
UniCUE (Ours) | 0.205 | 6.729 | 7.632 | 2.46 | 0.53 | 0.248 | 6.491 | 8.076 | 2.17 |
@@ 4. Experiment
4.1. Experimental Setting
Dataset. Existing CS datasets are limited to normal-hearing cuers and lack data from hearing-impaired individuals, hindering model generalization to the primary users of assistive communication systems. To bridge this gap, we construct a new dataset, the Unified-HI Corpus, which includes CS videos from 8 hearing-impaired and 6 normal-hearing cuers. This diverse composition significantly enriches variations in gesture styles, lip movements, and speech patterns. The expanded coverage introduces more realistic challenges and better reflects practical use cases, enabling models to capture cue-specific nuances essential for hearing-impaired users. A comparison with existing CS datasets is shown in Table 1, and further details on sentence coverage and phoneme distribution are included in Appendix Section 2.
Due to the noisy speech data from hearing-impaired cuers, we use CS data from 6 normal-hearing cuers for training. The data from normal-hearing cuers is split by sentence into training and test sets with a 95:5 ratio to ensure effective training and validation. Importantly, all CS data from the 8 hearing-impaired cuers are used in the test set, enabling a robust evaluation of model generalization to this group.
Architecture Details. The CSV2S pathway is entirely built upon the AudioLDM (Liu et al., 2023), including its VAE encoder-decoder, latent diffusion model, and vocoder components. For CSR, the Transformer in visual process, tokenizer, text-ViT, and text decoder are initialized from MBart (Liu et al., 2020b). Detailed training and inference configurations are provided in Appendix Section 1.
Evaluation Metrics. We evaluate the synthesized speech from three perspectives: linguistic accuracy, temporal synchronization, and speech quality. Linguistic accuracy is quantified by the Word Error Rate (WER) between the recognized text and ground truth. Temporal synchronization is assessed using SyncNet (Chung and Zisserman, 2016), reporting LSE-D (temporal distance) and LSE-C (confidence score). Speech quality is evaluated via STOI (Taal et al., 2010) for intelligibility and DNSMOS (Reddy et al., 2021) for naturalness.
Comparison Methods. We evaluate our UniCUE against: (1) CSV2S (Ours): direct speech synthesis without CSR assistance; (2) CSR (Ours): including pose-aware visual processor, text encoder and decoder, and semantic alignment pool; (3) CSR methods: CMML (Liu and Liu, 2023) and EcoCued (Liu et al., 2024); (4) V2S methods: Lip2Speech (Choi et al., 2023) and LipVoicer (Yemini et al., 2024).
@@ 4.2. Comparison with SOTA Methods
Quantitative Comparison. We compare our framework against SOTA methods, as summarized in Table 2. Our CSR model, empowered by the pose-aware visual processor and semantic alignment pool, achieves significantly lower WERs (0.186 for normal-hearing and 0.224 for hearing-impaired cuers), surpassing previous CSR methods. Building on this strong semantic understanding, UniCUE outperforms V2S methods across LSE-D, LSE-C, DNSMOS, and STOI metrics, demonstrating superior linguistic accuracy, temporal alignment, and speech quality.
Qualitative Comparison. Mel-spectrogram visualizations (Figure 4 in Appendix) further highlight the advantages of our method, showcasing improved temporal synchronization and clearer acoustic structures compared to others.
Method | Normal-hearing cuers | Hearing-impaired cuers | |||||||
WER | LSE-C | LSE-D | DNSMOS | STOI | WER | LSE-C | LSE-D | DNSMOS | |
GT | - | 7.274 | 7.314 | 2.79 | - | - | - | - | - |
CSR†† | 0.210 | 4.746 | 9.129 | 2.42 | 0.49 | 0.250 | 3.218 | 9.402 | 2.19 |
CSR‡ | 0.204 | 4.783 | 9.224 | 2.46 | 0.53 | 0.247 | 3.234 | 9.397 | 2.21 |
CSR | 0.186 | 4.874 | 9.125 | 2.53 | 0.57 | 0.224 | 3.342 | 9.315 | 2.29 |
CSV2S† | 0.398 | 6.158 | 8.122 | 2.21 | 0.40 | 0.398 | 5.821 | 8.582 | 1.96 |
CSV2S | 0.374 | 6.245 | 7.962 | 2.27 | 0.42 | 0.422 | 5.938 | 8.347 | 2.04 |
UniCUE†† | 0.239 | 6.637 | 7.724 | 2.30 | 0.44 | 0.267 | 6.419 | 8.163 | 2.08 |
UniCUE‡ | 0.231 | 6.641 | 7.716 | 2.33 | 0.46 | 0.276 | 6.421 | 8.159 | 2.10 |
UniCUE* | 0.226 | 6.613 | 7.731 | 2.37 | 0.48 | 0.271 | 6.410 | 8.167 | 2.12 |
UniCUE | 0.205 | 6.729 | 7.632 | 2.46 | 0.53 | 0.248 | 6.491 | 8.076 | 2.17 |
@@ 4.3. Ablation Studies
To verify the contribution of each component, we conduct ablation studies on both normal-hearing and hearing-impaired test data. Results are summarized in Table 3.
Unified Training Paradigm. Compared to direct CSV2S, UniCUE reduces WER by 45% (0.205 vs. 0.374) on normal-hearing cuers and 41% (0.248 vs. 0.422) on hearing-impaired cuers. These results highlight the benefit of leveraging fine-grained visual semantics from CSR to enhance CSV2S, alleviating the challenge of modeling complex multimodal correlations.
Visual Processor Design. Models that rely solely on raw video features struggle to capture fine-grained motion due to redundant and noisy visual information, resulting in suboptimal performance. By incorporating pose cues, our visual processor effectively captures cuer-specific dynamics, leading to significantly improved accuracy and robustness across diverse cuers. Semantic Alignment Mechanism. Disabling the Semantic Alignment Pool (SAP) degrades visual-semantic consistency, resulting in higher WERs for both CSR and UniCUE. This underscores the importance of the alignment in enforcing spatiotemporal coherence between visual cues and phonemic representations for accurate semantic modeling. The effectiveness of SAP is further validated by the t-SNE visualizations (Figure 5 in Appendix).
VisioPhonetic Adapter. Removing the VPA results in noticeable degradation in temporal alignment, demonstrating its crucial role in bridging the representation gap between CSR and CSV2S. By adaptively selecting and refining fine-grained spatialtemporal visual cues through learnable queries, the VPA enables more accurate and temporally coherent speech synthesis.
Impact of Hand Cues. Removing hand cues leads to substantial performance degradation, particularly for hearing-impaired users who often exhibit limited oral articulation and atypical lip shapes (see Appendix Table 1). The results highlight the complementary role of hand gestures in enhancing visual phonemic representations for CS.
Computational Efficiency. UniCUE achieves faster training convergence and 40% lower inference time than the combined pipeline (see details in Appendix Section 5).